# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v1

@@ Appendix B Notation and Basic Lemmas
Suppose $\widehat{Q}_{k,m}^{0}:=0$ and let $\widehat{Q}_{k,m}^{t+1}(s_{g},F_{z_{\Delta}},a_{g})=\widehat{\mathcal{T}}_{k,m}\widehat{Q}_{k,m}^{t}(s_{g},F_{z_{\Delta}},a_{g})$ for $t\in\mathbb{N}$. Denote the fixed-point of $\widehat{\mathcal{T}}_{k,m}$ by $\widehat{Q}^{\mathrm{est}}_{k,m}$ such that $\widehat{\mathcal{T}}_{k,m}\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},F_{z_{\Delta}},a_{g})=\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},F_{z_{\Delta}},a_{g})$.
Corollary B.11.1.
@@
Observe that by backpropagating results of the $\gamma$-contractive property for $T$ time steps:
$\|\widehat{{Q}}_{k}^{*}-\widehat{{Q}}_{k}^{T}\|_{\infty}\leq\gamma^{T}\cdot\|\widehat{Q}_{k}^{*}-\widehat{Q}_{k}^{0}\|_{\infty},\quad\quad\quad\|\widehat{{Q}}_{k,m}^{\mathrm{est}}-\widehat{{Q}}_{k,m}^{T}\|_{\infty}\leq\gamma^{T}\cdot\|\widehat{Q}_{k,m}^{\mathrm{est}}-\widehat{Q}^{0}_{k,m}\|_{\infty}$ (22)
@@
Further, noting that $\widehat{Q}_{k}^{0}=\widehat{Q}_{k,m}^{0}:=0$, $\|\widehat{Q}_{k}^{*}\|_{\infty}\leq\frac{\widetilde{r}}{1-\gamma}$, and $\|\widehat{Q}_{k,m}^{\mathrm{est}}\|_{\infty}\leq\frac{\widetilde{r}}{1-\gamma}$ from Lemma B.5:
$\|\widehat{Q}_{k}^{*}-\widehat{Q}_{k}^{T}\|_{\infty}\leq\gamma^{T}\frac{\widetilde{r}}{1-\gamma},\quad\quad\quad\|\widehat{Q}_{k,m}^{\mathrm{est}}-\widehat{Q}_{k,m}^{T}\|_{\infty}\leq\gamma^{T}\frac{\widetilde{r}}{1-\gamma}$ (23)
Remark B.12.
@@
Corollary B.11.1 characterizes the error decay between $\widehat{Q}_{k}^{T}$ and $\widehat{Q}_{k}^{*}$ and shows that it decays exponentially in the number of Bellman iterations by a $\gamma^{T}$ multiplicative factor.
@@
Furthermore, we characterize the maximal policies greedy policies obtained from $Q^{*},\widehat{Q}_{k}^{*}$, and $\widehat{Q}_{k,m}^{\mathrm{est}}$.
Definition B.13 (Optimal policy $\pi^{*}$).
@@
Definition B.14 (Optimal subsampled policy $\widehat{\pi}_{k}^{*}$).
The greedy policy from $\widehat{Q}_{k}^{*}$ is
$\widehat{\pi}_{k}^{*}(s_{g},s_{i},F_{s_{\Delta\setminus i}})\coloneq\arg\max_{(a_{g},a_{i},F_{a_{\Delta\setminus i}})\in\mathcal{A}_{g}\times\mathcal{A}_{l}\times\mu_{k-1}{(\mathcal{A}_{l})}}\widehat{Q}_{k}^{*}(s_{g},s_{i},F_{z_{\Delta\setminus i}},a_{i},a_{g}).$ (25)
Definition B.15 (Optimal empirically subsampled policy $\widehat{\pi}_{k,m}^{\mathrm{est}}$).
The greedy policy from $\widehat{Q}_{k,m}^{\mathrm{est}}$ is given by
$\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}}):=\arg\max_{(a_{g},a_{i},F_{a_{\Delta\setminus i}})\in\mathcal{A}_{g}\times\mathcal{A}_{l}\times\mu_{k-1}{(\mathcal{A}_{l})}}\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{z_{\Delta\setminus i}},a_{i},a_{g}).$ (26)
Figure 4 details the analytic flow on how we use the empirical adapted Bellman operator to perform value iteration on $\widehat{Q}_{k,m}$ to get $\widehat{Q}_{k,m}^{\mathrm{est}}$ which approximates $Q^{*}$.
Algorithm 4 gives a stable implementation of Algorithm 2 with learning rates $\{\eta_{t}\}_{t\in[T]}$. Algorithm 4 is provably numerical stable under fixed-point arithmetic (Anand et al., 2024). $\widehat{Q}_{k,m}^{t}$ in Algorithm 4 follows the $\gamma$-contractivity as in Lemma B.7, given an appropriately conditioned sequence of learning rates $\eta_{t}$:
@@
2:  Set $\widehat{Q}^{0}_{k,m}(s_{g},s_{1},F_{z_{\Delta\setminus 1}},a_{1},a_{g})=0$ for $(s_{g},s_{1},F_{z_{\Delta\setminus 1}},a_{1},a_{g})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}{(\mathcal{Z}_{l})}\times\mathcal{A}_{l}\times\mathcal{S}_{g}$.
3:  for $t=1$ to $T$ do
4:     for $(s_{g},s_{1},F_{z_{\Delta\setminus 1}})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}^{(\mathcal{Z}_{l})}$ do
5:        for $(a_{g},a_{1})\in\mathcal{A}_{g}\times\mathcal{A}_{l}$ do
6:           $\widehat{Q}^{t+1}_{k,m}(s_{g},s_{1},F_{z_{\Delta\setminus 1}},a_{g},a_{1})\leftarrow(1-\eta_{t})\widehat{Q}_{k,m}^{t}(s_{g},s_{1},F_{z_{\Delta\setminus 1}},a_{g},a_{1})+\eta_{t}\widehat{\mathcal{T}}_{k,m}\widehat{Q}^{t}_{k,m}(s_{g},s_{1},F_{z_{\Delta\setminus 1}},a_{g},a_{1})$
7:  Let the approximate policy be
$\widehat{\pi}_{k,m}^{T}(s_{g},s_{1},F_{s_{\Delta\setminus 1}})=\arg\max_{(a_{g},a_{1},a_{\Delta\setminus 1})\in\mathcal{A}_{g}\times\mathcal{A}_{l}\times\mu_{k-1}{(\mathcal{A}_{l})}}\widehat{Q}_{k,m}^{T}(s_{g},s_{1},F_{z_{\Delta\setminus 1}},a_{g},a_{1}).$
Theorem B.16.
@@
Furthermore, finite-time guarantees with the learning rate and sample complexity have been shown recently in Chen & Theja Maguluri (2022), which when adapted to our $\widehat{Q}_{k,m}$ framework in Algorithm 4 yields:
Theorem B.17 (Chen & Theja Maguluri (2022)).
For all $t\in[T]$ and $\epsilon>0$, if $\eta_{t}=(1-\gamma)^{4}\epsilon^{2}$ and $T=|\mathcal{S}_{l}||\mathcal{A}_{l}|k^{|\mathcal{S}_{l}||\mathcal{A}_{l}|}|\mathcal{S}_{g}||\mathcal{A}_{g}|/(1-\gamma)^{5}\epsilon^{2}$,
$\|\widehat{Q}_{k,m}^{T}-\widehat{Q}_{k,m}^{\mathrm{est}}\|\leq\epsilon.$
Definition B.18 (Star Graph $S_{n}$).
@@
$S_{n}$ captures the graph density notion by saturating the set of neighbors of the central node. Such settings find applications beyond reinforcement learning as well (Chaudhari et al., 2024; Anand & Qu, 2024; Li et al., 2019). The cardinality of the search space simplex for the optimal policy is exponential in $n$, so it cannot be naively modeled by an MDP: we need to exploit the symmetry of the local agents. This intuition allows our subsampling algorithm to run in polylogarithmic time (in $n$). Some works leverage an exponential decaying property that truncates the search space for policies over immediate neighborhoods of agents; however, this still relies on the assumption that the graph neighborhood for the agent is sparse (Qu et al., 2020a, b); however, $S_{n}$ is not locally sparse; hence, previous methods do not apply to this problem instance.
@@
Appendix C Proof Sketch
This section details an outline for the proof of Theorem 3.4, as well as some key ideas. At a high level, our SUBSAMPLE-MFQ framework recovers exact mean-field $Q$ learning and traditional value iteration when $k=n$ and as $m\to\infty$. Further, as $k\!\to\!n$, $\widehat{Q}_{k}^{*}$ should intuitively get closer to $Q^{*}$ from which the optimal policy is derived. Thus, the proof is divided into three major steps: firstly, we prove a Lipschitz continuity bound between $\widehat{Q}_{k}^{*}$ and $\widehat{Q}_{n}^{*}$ in terms of the total variation (TV) distance between $F_{z_{\Delta}}$ and $F_{z_{[n]}}$. Next, we bound the TV distance between $F_{z_{\Delta}}$ and $F_{z_{[n]}}$. Finally, we bound the value differences between ${\pi}_{k}^{\mathrm{est}}$ and $\pi^{*}$ by bounding $Q^{*}(s,\pi^{*}(s))-Q^{*}(s,{\pi}_{k}^{\mathrm{est}}(s))$ and then using the performance difference lemma from Kakade & Langford (2002).
Step 1: Lipschitz Continuity Bound. To compare $\widehat{Q}_{k}^{*}(s_{g},F_{s_{\Delta}},a_{g})$ with $Q^{*}(s,a_{g})$, we prove a Lipschitz continuity bound between $\widehat{Q}^{*}_{k}(s_{g},F_{s_{\Delta}},a_{g})$ and $\widehat{Q}^{*}_{k^{\prime}}(s_{g},F_{s_{\Delta^{\prime}}},a_{g})$ with respect to the TV distance measure between $s_{\Delta}\in\binom{s_{[n]}}{k}$ and $s_{\Delta^{\prime}}\in\binom{s_{[n]}}{k^{\prime}}$:
Theorem C.1 (Lipschitz continuity in $\widehat{Q}_{k}^{*}$).
@@ Appendix C Proof Sketch
$|\widehat{Q}^{*}_{k}(s_{g},F_{z_{\Delta}},a_{g})$ $-\widehat{Q}^{*}_{k^{\prime}}(s_{g},F_{z_{\Delta^{\prime}}},a_{g})|\leq{2(1-\gamma)^{-1}\|r_{l}(\cdot,\cdot)\|_{\infty}}\cdot\mathrm{TV}\left(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}}\right)$
We defer the proof of Theorem C.1 to Appendix D. See Fig. 4 for a comparison between the $\widehat{Q}_{k}^{*}$ learning and estimation process, and the exact ${Q}$-learning framework.
${\widehat{Q}^{0}_{k,m}(s_{g},F_{z_{\Delta}},a_{g})}$${Q^{*}(s_{g},s_{[n]},a_{g},a_{[n]})}$${\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},F_{z_{\Delta}},a_{g})}$${\widehat{Q}_{k}^{*}(s_{g},F_{z_{\Delta}},a_{g})}$${\widehat{Q}_{n}^{*}(s_{g},F_{z_{[n]}},a_{g})}$$\scriptstyle{\begin{subarray}{c}(1)\end{subarray}}$$\scriptstyle{\begin{subarray}{c}(2)\end{subarray}}$$\scriptstyle{\begin{subarray}{c}(3)\\ \approx\end{subarray}}$$\scriptstyle{\begin{subarray}{c}(4)\\ =\end{subarray}}$
Figure 4: Flow of the algorithm and relevant analyses in learning $Q^{*}$. Here, (1) follows by performing Algorithm 2 (SUBSAMPLE-MFQ: Learning) on $\widehat{Q}_{k,m}^{0}$. (2) follows from Lemma 3.2. (3) follows from the Lipschitz continuity and total variation distance bounds in Theorems C.1 and C.2. Finally, (4) follows from noting that $\widehat{Q}_{n}^{*}=Q^{*}$.
@@
We bound the TV distance between $F_{z_{\Delta}}$ and $F_{z_{[n]}}$, where $\Delta\!\in\!\mathcal{U}\binom{[n]}{k}$. This task is equivalent to bounding the discrepancy between the empirical distribution and the distribution of the underlying finite population. When each $i\!\in\!\Delta$ is uniformly sampled without replacement, we use Lemma E.3 from Anand & Qu (2024) which generalizes the Dvoretzky-Kiefer-Wolfowitz (DKW) concentration inequality for empirical distribution functions. Using this, we show:
Theorem C.2.
@@
Then, by Theorem C.2 and the definition of TV distance from Section 2, we have that for $\delta\in(0,1]$, with probability at least $1-\delta$,
@@
Step 3: Performance Difference Lemma to Complete the Proof. As a consequence of the prior two steps and Lemma 3.2, $Q^{*}(s,a^{\prime})$ and $\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},F_{z_{\Delta}},a_{g}^{\prime})$ become similar as $k\to n$. We further prove that the value generated by the policies $\pi^{*}$ and ${\pi}_{k}^{\mathrm{est}}$ must also be very close (where the residue shrinks as $k\to n$). We then use the well-known performance difference lemma (Kakade & Langford, 2002) which we restate in Appendix F.1. A crucial theorem needed to use the performance difference lemma is a bound on $Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-Q^{*}(s^{\prime},\widehat{\pi}_{k}^{\mathrm{est}}(s_{g}^{\prime},F_{s_{\Delta}^{\prime}}))$. Therefore, we formulate and prove Theorem C.3 which yields a probabilistic bound on this difference, where the randomness is over the choice of $\Delta\in\binom{[n]}{k}$:
Theorem C.3.
@@
$Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-Q^{*}(s^{\prime},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g}^{\prime},F_{s_{\Delta}^{\prime}}))\leq\frac{2\|r_{l}(\cdot,\cdot)\|_{\infty}}{1-\gamma}\sqrt{\frac{n-k+1}{2nk}\ln\left(\frac{2|\mathcal{Z}_{l}|}{\delta}\right)}+2\epsilon_{k,m}.$
We defer the proof of Theorem C.3 and finding optimal value of $\delta$ to Lemma F.7 in the Appendix. Using Theorem C.3 and the performance difference lemma leads to Theorem 3.4.
Appendix D Proof of Lipschitz-continuity Bound
This section proves a Lipschitz-continuity bound between $\widehat{Q}_{k}^{*}$ and $Q^{*}$ and includes a framework to compare $\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}^{*}_{k}(s_{g},s_{\Delta},a_{g})$ and $Q^{*}(s,a_{g})$ in Lemma D.11.
Let $z_{i}:=(s_{i},a_{i})\in\mathcal{Z}_{l}\coloneq\mathcal{S}_{l}\times\mathcal{A}_{l}$ and $z_{\Delta}=\{z_{i}:i\in\Delta\}\in\mathcal{Z}_{l}^{k}$. For $z_{i}=(s_{i},a_{i})$, let $z_{i}(s)=s_{i}$, and $z_{i}(a)=a_{i}$. With abuse of notation, note that $\widehat{Q}_{k}^{T}(s_{g},a_{g},s_{\Delta},a_{\Delta})$ is equal to $\widehat{Q}_{k}^{T}(s_{g},a_{g},z_{\Delta})$.
The following definitions will be relevant to the proof of Theorem D.3.
Definition D.1 (Empirical Distribution Function).
For all $z\in\mathcal{Z}_{l}^{|\Delta|}$ and $s^{\prime},a^{\prime}\in\mathcal{S}_{l}\times\mathcal{A}_{l}$, where $\Delta\subseteq[n]$,

### v3

@@ Appendix C Notation and Basic Lemmas
Let $\hat{Q}_{k,m}^{0}:=0$ and $\hat{Q}_{k,m}^{t+1}(s_{g},F_{z_{\Delta}},a_{g})=\hat{\mathcal{T}}_{k,m}\hat{Q}_{k,m}^{t}(s_{g},F_{z_{\Delta}},a_{g})$ for $t\in\mathbb{N}$. Then, the Banach-Cacciopoli fixed-point of the adapted Bellman operator $\hat{\mathcal{T}}_{k,m}$ is $\hat{Q}^{\mathrm{est}}_{k,m}$ such that $\hat{\mathcal{T}}_{k,m}\hat{Q}^{\mathrm{est}}_{k,m}(s_{g},F_{z_{\Delta}},a_{g})=\hat{Q}^{\mathrm{est}}_{k,m}(s_{g},F_{z_{\Delta}},a_{g})$.
Corollary C.11.
@@
Observe that by recursively using the $\gamma$-contractive property for $T$ time steps:
$\|\hat{{Q}}_{k}^{*}-\hat{{Q}}_{k}^{T}\|_{\infty}\leq\gamma^{T}\cdot\|\hat{Q}_{k}^{*}-\hat{Q}_{k}^{0}\|_{\infty}$ (17)
$\|\hat{{Q}}_{k,m}^{\mathrm{est}}-\hat{{Q}}_{k,m}^{T}\|_{\infty}\leq\gamma^{T}\cdot\|\hat{Q}_{k,m}^{\mathrm{est}}-\hat{Q}^{0}_{k,m}\|_{\infty}$ (18)
@@
Further, noting that $\hat{Q}_{k}^{0}=\hat{Q}_{k,m}^{0}:=0$, $\|\hat{Q}_{k}^{*}\|_{\infty}\leq\frac{\tilde{r}}{1-\gamma}$, and $\|\hat{Q}_{k,m}^{\mathrm{est}}\|_{\infty}\leq\frac{\tilde{r}}{1-\gamma}$ from Lemma C.4:
$\|\hat{Q}_{k}^{*}-\hat{Q}_{k}^{T}\|_{\infty}\leq\gamma^{T}\frac{\tilde{r}}{1-\gamma},$ (19)
$\|\hat{Q}_{k,m}^{\mathrm{est}}-\hat{Q}_{k,m}^{T}\|_{\infty}\leq\gamma^{T}\frac{\tilde{r}}{1-\gamma}$ (20)
Remark C.12.
@@
Corollary C.11 characterizes the error decay between $\hat{Q}_{k}^{T}$ and $\hat{Q}_{k}^{*}$ and shows that it decays exponentially in the number of Bellman iterations by a $\gamma^{T}$ multiplicative factor.
@@
Furthermore, we characterize the maximal policies greedy policies obtained from $Q^{*},\hat{Q}_{k}^{*}$, and $\hat{Q}_{k,m}^{\mathrm{est}}$.
Definition C.13 (Optimal policy $\pi^{*}$).
@@
Definition C.14 (Optimal subsampled policy $\hat{\pi}_{k}^{*}$).
The greedy policy from $\hat{Q}_{k}^{*}$ is
$\hat{\pi}_{k}^{*}(s_{g},s_{i},F_{s_{\Delta\setminus i}})\coloneq\mathop{\operatorname*{arg\,max}}_{(a_{g},a_{i},F_{a_{\Delta\setminus i}})\in\mathcal{A}_{g}\times\mathcal{A}_{l}\times\mu_{k-1}{(\mathcal{A}_{l})}}\hat{Q}_{k}^{*}(s_{g},s_{i},F_{z_{\Delta\setminus i}},a_{i},a_{g}).$ (22)
Definition C.15 (Optimal empirically subsampled policy $\hat{\pi}_{k,m}^{\mathrm{est}}$).
The greedy policy from $\hat{Q}_{k,m}^{\mathrm{est}}$ is given by
$\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}}):=\mathop{\operatorname*{arg\,max}}_{(a_{g},a_{i},F_{a_{\Delta\setminus i}})\in\mathcal{A}_{g}\times\mathcal{A}_{l}\times\mu_{k-1}{(\mathcal{A}_{l})}}\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{z_{\Delta\setminus i}},a_{i},a_{g}).$ (23)
Figure 4 details the analytic flow on how we use the empirical adapted Bellman operator to perform value iteration on $\hat{Q}_{k,m}$ to get $\hat{Q}_{k,m}^{\mathrm{est}}$ which approximates $Q^{*}$.
Algorithm 3 gives a stable implementation of Algorithm 1 with learning rates $\{\eta_{t}\}_{t\in[T]}$. Algorithm 3 is provably numerical stable under fixed-point arithmetic [Anand et al., 2024, 2025]. $\hat{Q}_{k,m}^{t}$ in Algorithm 3 is $\gamma$-contractive as in Lemma C.6, given an appropriately conditioned sequence of learning rates $\eta_{t}$:
@@
2:  for $(s_{g},s_{1},F_{z_{\Delta\setminus 1}})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}{(\mathcal{Z}_{l})}$ do
3:     for $(a_{g},a_{1})\in\mathcal{A}_{g}\times\mathcal{A}_{l}$ do
4:        Set $\hat{Q}^{0}_{k,m}(s_{g},s_{1},F_{z_{\Delta\setminus 1}},a_{1},a_{g})=0$
5:  for $t=1$ to $T$ do
6:     for $(s_{g},s_{1},F_{z_{\Delta\setminus 1}})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}{(\mathcal{Z}_{l})}$ do
7:        for $(a_{g},a_{1})\in\mathcal{A}_{g}\times\mathcal{A}_{l}$ do
8:
$\hat{Q}^{t+1}_{k,m}(s_{g},s_{1},F_{z_{\Delta\setminus 1}},a_{g},a_{1})$ $=(1-\eta_{t})\hat{Q}_{k,m}^{t}(s_{g},s_{1},F_{z_{\Delta\setminus 1}},a_{g},a_{1})$
$\quad\quad\quad\quad+\!\eta_{t}\hat{\mathcal{T}}_{k,m}\hat{Q}^{t}_{k,m}(s_{g},s_{1},F_{z_{\Delta\setminus 1}},a_{g},a_{1})$
9:  Let the approximate policy be
$\hat{\pi}_{k,m}^{T}(s_{g},s_{1},F_{s_{\Delta\setminus 1}})=\mathop{\operatorname*{arg\,max}}_{(a_{g},a_{1},a_{\Delta\setminus 1})\in\mathcal{A}_{g}\times\mathcal{A}_{l}\times\mu_{k-1}{(\mathcal{A}_{l})}}\hat{Q}_{k,m}^{T}(s_{g},s_{1},F_{z_{\Delta\setminus 1}},a_{g},a_{1}).$
Theorem C.16.
@@
Furthermore, finite-time guarantees with the learning rate and sample complexity have been shown in Chen and Maguluri [2022b], which when adapted to our $\hat{Q}_{k,m}$ framework in Algorithm 3 yields:
Theorem C.17 (Chen and Maguluri [2022b]).
For all $t=\{1,\dots,T\}$ and for any $\epsilon>0$, if the learning rate sequence $\eta_{t}$ satisfies $\eta_{t}=(1-\gamma)^{4}\epsilon^{2}$ and $T=\frac{|\mathcal{S}_{l}||\mathcal{A}_{l}|k^{|\mathcal{S}_{l}||\mathcal{A}_{l}|}|\mathcal{S}_{g}||\mathcal{A}_{g}|}{(1-\gamma)^{5}}\epsilon^{2}$, then
$\|\hat{Q}_{k,m}^{T}-\hat{Q}_{k,m}^{\mathrm{est}}\|\leq\epsilon.$
Definition C.18 (Star Graph $S_{n}$).
@@
$S_{n}$ captures the graph density notion by saturating the set of neighbors of the central node. Such settings find applications beyond RL as well [Chaudhari et al., 2024, Anand and Qu, 2024, Li et al., 2019]. The cardinality of the search space simplex for the optimal policy is exponential in $n$, so it cannot be naively modeled by an MDP: we need to exploit the symmetry of the local agents. This intuition allows our subsampling algorithm to run in polylogarithmic time (in $n$). Some works leverage an exponential decaying property that truncates the search space for policies over immediate neighborhoods of agents; however, this still relies on the assumption that the graph neighborhood for the agent is sparse [Qu et al., 2020a, b]; however, $S_{n}$ is not locally sparse; hence, previous methods do not apply to this problem instance.
@@
Finally, we argue why mean-field value iteration faithfully performs the same update as value iteration.
Lemma C.19 (Equivalence of Mean-Field Value Iteration and Value Iteration).
Since the local agents $1,\dots,n$ are all homogeneous in their state/action spaces, the $\hat{Q}_{k}$-function only depends on them through their empirical distribution $F_{z_{\Delta}}$, proving the lemma. Therefore, for the remainder of the paper, we will use $\hat{Q}_{k}(s_{g},F_{z_{\Delta}},a_{g})\coloneq\hat{Q}_{k}(s_{g},s_{i},F_{z_{\Delta\setminus i}},a_{g},a_{i})\coloneq\hat{Q}_{k}(s_{g},s_{\Delta},a_{g},a_{\Delta})$ interchangeably, unless making a remark about the computational complexity of learning each function.
Appendix D Proof Sketch
This section details an outline for the proof of Theorem 4.4, as well as some key ideas. At a high level, our SUBSAMPLE-MFQ framework recovers exact mean-field $Q$ learning and traditional value iteration when $k=n$ and as $m\to\infty$. Further, as $k\!\to\!n$, $\hat{Q}_{k}^{*}$ should intuitively get closer to $Q^{*}$ from which the optimal policy is derived. Thus, the proof is divided into three major steps: firstly, we prove a Lipschitz continuity bound between $\hat{Q}_{k}^{*}$ and $\hat{Q}_{n}^{*}$ in terms of the total variation (TV) distance between $F_{z_{\Delta}}$ and $F_{z_{[n]}}$. Next, we bound the TV distance between $F_{z_{\Delta}}$ and $F_{z_{[n]}}$. Finally, we bound the value differences between ${\pi}_{k}^{\mathrm{est}}$ and $\pi^{*}$ by bounding $Q^{*}(s,\pi^{*}(s))-Q^{*}(s,{\pi}_{k}^{\mathrm{est}}(s))$ and then using the performance difference lemma from Kakade and Langford [2002].
Step 1: Lipschitz Continuity Bound. To compare $\hat{Q}_{k}^{*}(s_{g},F_{s_{\Delta}},a_{g})$ with $Q^{*}(s,a_{g})$, we prove a Lipschitz continuity bound between $\hat{Q}^{*}_{k}(s_{g},F_{s_{\Delta}},a_{g})$ and $\hat{Q}^{*}_{k^{\prime}}(s_{g},F_{s_{\Delta^{\prime}}},a_{g})$ with respect to the TV distance measure between $s_{\Delta}\in\binom{s_{[n]}}{k}$ and $s_{\Delta^{\prime}}\in\binom{s_{[n]}}{k^{\prime}}$:
Theorem D.1 (Lipschitz continuity in $\hat{Q}_{k}^{*}$).
@@ Appendix D Proof Sketch
$|\hat{Q}^{*}_{k}(s_{g},F_{z_{\Delta}},a_{g})$ $-\hat{Q}^{*}_{k^{\prime}}(s_{g},F_{z_{\Delta^{\prime}}},a_{g})|\leq\frac{2}{1-\gamma}\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}\left(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}}\right)$
We defer the proof of Theorem D.1 to Appendix E. See Figure 4 for a comparison between the $\hat{Q}_{k}^{*}$ learning and estimation process, and the exact ${Q}$-learning framework.
${\hat{Q}^{0}_{k,m}(s_{g},F_{z_{\Delta}},a_{g})}$${Q^{*}(s_{g},s_{[n]},a_{g},a_{[n]})}$${\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},F_{z_{\Delta}},a_{g})}$${\hat{Q}_{k}^{*}(s_{g},F_{z_{\Delta}},a_{g})}$${\hat{Q}_{n}^{*}(s_{g},F_{z_{[n]}},a_{g})}$$\scriptstyle{\begin{subarray}{c}(1)\end{subarray}}$$\scriptstyle{\begin{subarray}{c}(2)\end{subarray}}$$\scriptstyle{\begin{subarray}{c}(3)\\ \approx\end{subarray}}$$\scriptstyle{\begin{subarray}{c}(4)\\ =\end{subarray}}$
Figure 4: Flow of the algorithm and relevant analyses in learning $Q^{*}$. Here, (1) follows by performing Algorithm 1 (SUBSAMPLE-MFQ: Learning) on $\hat{Q}_{k,m}^{0}$. (2) follows from Lemma 4.3. (3) follows from the Lipschitz continuity and total variation distance bounds in Theorems D.1 and D.2. Finally, (4) follows from noting that $\hat{Q}_{n}^{*}=Q^{*}$.
@@
We bound the TV distance between $F_{z_{\Delta}}$ and $F_{z_{[n]}}$, where $\Delta\!\in\!\mathcal{U}\binom{[n]}{k}$. This task is equivalent to bounding the discrepancy between the empirical distribution and the distribution of the underlying finite population. When each $i\!\in\!\Delta$ is uniformly sampled without replacement, we use Lemma F.3 from Anand and Qu [2024] which generalizes the Dvoretzky-Kiefer-Wolfowitz (DKW) concentration inequality for empirical distribution functions.
Using this, we show:
Theorem D.2.
@@
Then, by Theorem D.2 and the definition of total variation distance from Section 2, we have that for $\delta\in(0,1]$, with probability at least $1-\delta$,
@@
Step 3: Performance Difference Lemma to Complete the Proof. As a consequence of the prior two steps and Lemma 4.3, $Q^{*}(s,a^{\prime})$ and $\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},F_{z_{\Delta}},a_{g}^{\prime})$ become similar as $k\to n$. We further prove that the value generated by the policies $\pi^{*}$ and ${\pi}_{k}^{\mathrm{est}}$ must also be very close (where the residue shrinks as $k\to n$). We then use the well-known performance difference lemma [Kakade and Langford, 2002] which we restate in Appendix G.1. A crucial theorem needed to use the performance difference lemma is a bound on $Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-Q^{*}(s^{\prime},\hat{\pi}_{k}^{\mathrm{est}}(s_{g}^{\prime},F_{s_{\Delta}^{\prime}}))$.
Therefore, we formulate and prove Theorem D.3 which yields a probabilistic bound on this difference, where the randomness is over the choice of $\Delta\in\binom{[n]}{k}$:
Theorem D.3.
@@
$Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-Q^{*}(s^{\prime},\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g}^{\prime},F_{s_{\Delta}^{\prime}}))\leq\frac{2\|r_{l}(\cdot,\cdot)\|_{\infty}}{1-\gamma}\sqrt{\frac{n-k+1}{2nk}\ln\left(\frac{2|\mathcal{Z}_{l}|}{\delta}\right)}+2\epsilon_{k,m}.$
We defer the proof of Theorem D.3 and finding optimal value of $\delta$ to Lemma G.8 in the Appendix. Using Theorem D.3 and the performance difference lemma leads to Theorem 4.4.
Appendix E Proof of Lipschitz-continuity Bound
This section proves a Lipschitz-continuity bound between $\hat{Q}_{k}^{*}$ and $Q^{*}$ and includes a framework to compare $\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\hat{Q}^{*}_{k}(s_{g},s_{\Delta},a_{g})$ and $Q^{*}(s,a_{g})$ in Lemma E.12.
Let $z_{i}:=(s_{i},a_{i})\in\mathcal{Z}_{l}\coloneq\mathcal{S}_{l}\times\mathcal{A}_{l}$ and $z_{\Delta}=\{z_{i}:i\in\Delta\}\in\mathcal{Z}_{l}^{k}$. For $z_{i}=(s_{i},a_{i})$, let $z_{i}(s)=s_{i}$, and $z_{i}(a)=a_{i}$. With abuse of notation, note that $\hat{Q}_{k}^{T}(s_{g},a_{g},s_{\Delta},a_{\Delta})$ is equal to $\hat{Q}_{k}^{T}(s_{g},a_{g},z_{\Delta})$.
The following definitions will be relevant to the proof of Theorem E.3.
Definition E.1 (Empirical Distribution Function).
For all $z\in\mathcal{Z}_{l}^{|\Delta|}$ and $(s^{\prime},a^{\prime})\in\mathcal{S}_{l}\times\mathcal{A}_{l}$, where $\Delta\subseteq[n]$,